OpenStudy (anonymous):

what is the solution of the following system? 4x-y=15 3x+2y=-8 (-1,-19) (2,-7) <--- is this the right answer? (4,-5) (5,5)

OpenStudy (anonymous):

So your system is \[4x-y=15\] \[3x+2y=-8\] scaling the first one up by 2, allows you to cancel out the y terms leaving you with \[11x=22\] So, x=2...which only is true in one of the choices...the one you picked :)

OpenStudy (texaschic101):

4x - y = 15 --->(2) 3x + 2y = -8 -------------- 8x - 2y = 30 (result of multiplying by 2) 3x + 2y = -8 --------------add 11x = 22 x = 2 now sub 2 in for x in either equation to find y 4x - y = 15 4(2) - y = 15 8 - y = 15 8 - 15 = y - 7 = y x = 2 and y = -7......you are correct